OpenStudy (anonymous):

Help! Write an exponential funct. y=ab^2 for a graph that includes (-1, 2.8),( 0,7) ? a. y=3.5(5)^x b. y=2.5(7)^x c. y=7(2.5)^x d. y=5(3.5)^x

OpenStudy (lucaz):

substitute x by the given x values and see if they give the corresponding y value

OpenStudy (lucaz):

fisrt equation y=3.5(5)^x, from the 1st point f(-1)=2.8 y=3.5(5)^-1=0.7, it doesnt work.. same process for the other equations (try with the 3rd equation)